OpenStudy (anonymous):

Simplify this expression: (3k)^0 k^-3/k^6 h^-9

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Is this it? \(\large \dfrac{(3k)^0 k^{-3}}{k^6 h^{-9}}\)

OpenStudy (triciaal):

are you going to answer?

OpenStudy (triciaal):

anything to the power zero is one to multiply numbers or variables with the same base add the exponents to divide numbers or variables with the same base subtract the exponents

OpenStudy (anonymous):

mathstudent55 That is right.

OpenStudy (michele_laino):

hint: the zero-th power of any quantity which is different from zero, is equal to \(1\), so, what is: \[\huge {\left( {3k} \right)^0} = ...?\]

OpenStudy (tkhunny):

...unless \(k = 0\)

OpenStudy (mathmale):

Zeke, if helpers need to ask you for clarification regarding your meaning, that's best taken as an indication that you should be drawing your question (in the Draw utility, below) or using Equation Editor. As a last resort, if you don't want to go through that trouble, use parentheses to indicate which operations are to be done first.

OpenStudy (mathmale):

Zeke? still there?